OpenStudy (anonymous):

Assume that x and y are positive numbers. Simplify by writing as one logarithm. (1/5)log(x^5) - (3/4)log(y^4) + 2log(radical xy) please explain

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 } \log (x ^{5}) - \frac{ 3 }{ 4 } \log (y ^{4}) + 2 \log ( \sqrt{xy})\]

OpenStudy (anonymous):

Can someone help me?

OpenStudy (anonymous):

@Loser66 can you help me with this question?

OpenStudy (loser66):

\(alog b = log b^a\) do you know that?

OpenStudy (anonymous):

yes i did know that

OpenStudy (loser66):

so, apply to the first term \(\dfrac{1}{5}log x^5\) what do you get?

OpenStudy (anonymous):

\[\log (x ^{5})^{\frac{ 1 }{ 5 }}\]

OpenStudy (anonymous):

right?

OpenStudy (loser66):

yes, and simplify more, please

OpenStudy (loser66):